OpenStudy (anonymous):

A circle has an area of 153.86 square inches. Using 3.14 for pi, find the radius and circumference of the circle. Round your answer to the nearest hundredth when necessary.

OpenStudy (johnweldon1993):

We know that Area = pi * r² right? so what do we need from this? we need the Radius because the Circumference is 2 * pi * r so Lets take the area that we know...153.86 and plug it into the equation we know Area = pi * r² 153.86 = pi * r² Now we know that pi = 3.14 from your question....so lets divide both sides by 3.14...this will isolate the r² 153.86 ------- = 49 3.14 So now we have 49 = r² Now we want 'r'...not r² ...so lets take the square root of both sides r = √49 r = ???

OpenStudy (anonymous):

@johnweldon1993 7 ?

OpenStudy (johnweldon1993):

Sorry about the late reply...yes 7 would be the radius... Now to find the Circumference....remember Circumference = 2 * pi * r We know pi = 3.14 via your question...and now we know that 'r' = 7...so Circumference = 2(3.14)(7) Circumference = ???